# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v1

@@ 7 Conclusions and Discussion
This paper addresses a critical roadblock in time series forecasting: information insufficiency. We introduced Text-Guided Time Series Forecasting (TGTSF), a new approach that integrates textual data to enrich the models with external information and system knowledge. We developed and released four TGTSF datasets, each crafted to validate different aspects of the task and model. Our straightforward yet effective TGForecaster model demonstrates that textual guidance can significantly enhance time series modeling by mitigating the average predictions typically resulting from information scarcity.
@@ Appendix A Data and Code Availability
Given the extensive size of the TGTSF dataset, an immediate anonymous release may not be feasible. Therefore, we plan to make the full dataset publicly available after the review period. Additionally, the script used to create the toy dataset will be released alongside our code base. You can access the anonymous code repository at: https://anonymous.4open.science/r/TGTSF_review-6E51.
@@ Appendix C Visualization Weather-Captioned Dataset
Appendix E Comparison with More Baselines on Electricity-Captioned
@@
Appendix F Full Result on Steam-100 Dataset
@@
Appendix G Toy Dataset Details
G.1 Data source
@@
G.2 Dataset Structure
@@
Appendix H Electricity-Caption Details
H.1 Data source
@@
H.2 Dataset Structure
@@
Appendix I Weather-Caption Details
I.1 Data source
We get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic.
I.2 Prompt
@@ I.4 Dataset Structure
Appendix J Steam Details
J.1 Data source
We directly crawl the event data on Steam News of each Game. As for the online gamers, we use a randomly picked publicly available database that provide the downloading of csv file containing online gamer number.
Appendix K Implementation Details and Hyper-Parameters
@@
We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. In most of the experiments, we simply use a patch length of 6 and stride of 3.

### v2

@@ 7 Conclusions and Discussion
This paper addresses a critical roadblock in time series forecasting: information insufficiency. We introduced Text-Guided Time Series Forecasting (TGTSF), a new approach that integrates textual cues to enrich the models with external information and system knowledge. We developed and released four TGTSF datasets, each crafted to validate different aspects of the task and model. Our straightforward yet effective TGForecaster model demonstrates that textual guidance can significantly enhance time series modeling by mitigating the average predictions typically resulting from information scarcity.
@@ Appendix A Data and Code Availability
Given the extensive size of the TGTSF dataset, especially the ones with pre-embeddings, an immediate anonymous release may not be feasible. Therefore, we plan to make the full datasets publicly available ASAP.
Additionally, the script used to create the toy dataset will be released alongside our code base. You can access the code repository at: https://github.com/VEWOXIC/TGTSF.
We will update the github repo here later and will update the dataset in it later.
@@ Appendix C Visualization Weather-Captioned Dataset
Appendix E Attention Map Visualization on Weather-Captioned
We further visualize two cross-attention blocks to further investigate the TGForecaster. You are strongly advised to check the Tab. 1, Appendix K.3 and Fig. 5 while reading this part.
Figure 7 illustrates the attention map of the "text-guided channel independent" cross-attention block in the text encoder across three layers. In the first layer, attention is predominantly focused on the first sentence, which specifies the month and time. This sentence is crucial as it provides temporal context that significantly impacts the prediction of both daily and annual periodicity. While other sentences receive moderate attention, the sixth sentence, which describes atmospheric pressure as detailed in Table 1, consistently receives no attention across all channels.
In the second layer, however, there is a notable shift in attention dynamics. All channels, particularly channel 0, show intense focus on the sixth sentence. According to the channel definitions in Appendix K.3, channel 0 directly corresponds to atmospheric pressure. Channels 10 and 20, which are related to air density and CO2 concentration respectively—factors closely associated with pressure—also display relatively high attention scores. This suggests that the TGForecaster is capable of discerning the underlying relationships among the channels.
The separation of attention focus between the first and second layers suggests that the influence of atmospheric pressure on the model’s predictions is independent of time. In the third layer, a diversity of attention patterns emerges; channel 0 focuses exclusively on the sixth sentence, while other channels predominantly attend to the first sentence.
Since we take the output of previous layer as query and input news embeddings as key and value, the information lies in the news are progressively added to the channel embeddings. Thus, the model can focus on different perspective in separate cross attention layers.
(a) Layer1
(b) Layer2
(c) Layer3
Figure 7: Attention map of the "text guided channel independent" cross attention block on weather-captioned dataset, on on the 15000th test sample of Weather-Caption-Medium dataset. We use three cross attention block. The vertical axis stand for channels and horizon stand for the 7 sentences of the weather report summary.
Figure 8 presents the attention map of the modality mixer layer cross attention block in the weather-captioned dataset. The map, averaged across three cross attention layers, illustrates distinct patterns of attention for each channel. This diversity underscores the TGForecaster’s ability to adaptively extract time series embeddings tailored to the unique distribution characteristics of each channel, facilitated by textual inputs.
Notably, the channels for SWDR, PAR, and max.PAR display clear periodic patterns in their attention maps, aligning with observations from waveform visualizations. These patterns suggest that the TGForecaster effectively captures and utilizes periodic information from these environmental variables.
Furthermore, the channels labeled rain and raining show a particularly interesting behavior; they assign significantly higher attention scores to the exact time periods of rainfall within the look-back window. This behavior indicates that the TGForecaster is adept at identifying and prioritizing crucial temporal events specific to each channel, further enhancing its forecasting accuracy by focusing on relevant patterns where needed. This level of detail in attention allocation demonstrates the model’s capability to integrate contextual cues from textual data and further guide the time series forecasting.
Figure 8: Attention map of the modality mixer layer cross attention block on weather-captioned dataset, onon the 15000th test sample of Weather-Caption-Medium dataset. The attention map is averaged across three cross attention layers. We plot the attention map for each channel. The vertical axis stand for output time series patches and the horizon stand for the input time series patches embedding from PatchTST backbone.
Appendix F Comparison with More Baselines on Electricity-Captioned
@@
Appendix G Full Result on Steam-100 Dataset
@@
Appendix H Error Bar & Critical Difference Diagram
We run the experiments on Toy and Electricity for five times with different randomly chosen random seeds. And Weather-Medium for three times because of the large amount of data can result in very long training time on our devices. We report the mean and standard deviation as follows with comparison with FITS, the most stable model.
As Tab. 9 indicate, TGForecaster shows stable performance across the benchmark. Even with extreme condition, it still maintains superior performance. It worth note that, we thought the relative large variance on weather dataset is caused by the different combination of the text description. But the FITS also shows large variance on this dataset which indicate it is hard to converge on this dataset.
We generate the critical difference plot on our result of four datasets (toy, Electricity, Weather-Medium, Weather-Large) with the default alpha as 0.05 as shown in Fig. 9. TGForecaster’s placement at the top of the critical difference plot, without intersecting with other lines, demonstrates its consistent and superior performance in terms of MSE compared to the other models. It indicates that with the help of external textual information, TGForecaster can handle complicated datasets.
Figure 9: The Critical Difference Plot on the TGForecaster and other baselines with alpha=0.05.
Table 9: The mean and std of TGForecaster on the three dataset in metrics of MSE.
Datasets TGForecaster FITS
Toy 0.027±0.001 0.883±0.000
Electricity 0.193±0.004 0.203±0.001
Weather-Medium 0.281±0.008 0.430±0.011
Appendix I Toy Dataset Details
I.1 Data source
@@
We will publish this dataset with CC BY-NC-SA 4.0 licence.
I.2 Dataset Structure
@@
Appendix J Electricity-Caption Details
J.1 Data source
@@
We will publish this dataset with CC BY-NC-SA 4.0 licence.
J.2 Dataset Structure
@@
Appendix K Weather-Caption Details
K.1 Data source
We get the weather time series data from: https://www.bgc-jena.mpg.de/wetter/ and weather report from https://www.timeanddate.com/weather/germany/jena/historic. We will publish this dataset with CC BY-NC-SA 4.0 licence since the data source forbids commercial use.
K.2 Prompt
@@ 5. Humidity: you should convert the humidity to: (<30%) -> Very Dry, (30-50%) -> Dry,
Appendix L Steam Details
L.1 Data source
We are not directly publishing this dataset because of the intellectual property restrictions.
For those who are interested in reproducing the dataset: We directly crawl all the event data on Steam News of each Game. We may release the script for this later.
As for the online players, we use a randomly picked publicly available database that provide the downloading of csv file containing online gamer number. We will not release any data or tools to get these data.
Appendix M Implementation Details and Hyper-Parameters
@@
For electricity dataset, we directly report the result from the original paper. For weather dataset, we uses the exact set of hyper-parameter for the original weather datasets provided by each baseline model.
In most of the experiments, we simply use a patch length of 6 and stride of 3. For Toy dataset, we use patch length of 16 and stride of 8.
We follow the previous works, split all the dataset by 7:1:2 for training, validation and testing.
@@
Further detailed hyperparameter settings are provided in the training scripts in our codebase. We did not perform comprehensive hyper-parameter tuning because of the constraint of compute power. Thus, we may report a sub-optimal result of TGForecaster.